# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

### v1

@@ 6.4. Hub-Highway Nodes Allow Queries to Traverse the Graph Faster
where $|B_{i}|$ is the bin size. We average this value over all queries and fix the bin size to $30$. Figure 10 shows results for Gist, GloVe, Microsoft SpaceV and the Yandex Deep benchmark datasets. We observe that earlier in the search procedure, queries tend to concentrate in the highway structures as shown by the percentage of hub nodes visited in the earlier bins. Since highway nodes are able to connect disparate regions in the graph, this result confirms that they propel queries to navigate the similarity search graph faster as they are visited the most earlier in the search. Worth noting is that for GloVe, as shown in Figure 10(b), the percentage of hub nodes visited is much lower across all the bins. This is consistent with prior findings that show that the cosine similarity exhibit anti-hub properties as it is not affected by the magnitude of the vectors compared with the $l_{2}$ distance.
@@
Figure 10. Highway nodes allow queries to traverse the graph faster.
@@ 7. Conclusion
Approximate near neighbor search has become an increasingly crucial computational workload in recent years with the seminal Hierarchical Navigable Small World (HNSW) algorithm continuing to garner significant interest and adoption from practitioners. Given the impact of HNSW on both the research community and production systems in industry, a deep understanding of the algorithm’s inner workings would be of immediate value.
To that end, this paper presents the first comprehensive study on the utility of the hierarchical component of HNSW over numerous large-scale datasets and performance metrics. Ultimately, we find that the hierarchy of HNSW provides no clear benefit on high-dimensional datasets and can be removed without any discernible loss in performance while providing memory savings.
@@
As part of our benchmarking study and analysis, we also release a high-performance flat navigable small world implementation, called FlatNav, that we hope will serve as an additional resource to both researchers and industry practitioners. We also hope that our work can inspire future directions in algorithm design for ANN search, such as in better strategies for exploiting the hub highway structure we observed or even designing new algorithms that do manage to leverage the idea of hierarchy effectively. In the short term, we believe that our results provide immediate implications for practitioners seeking to save memory or simplify their vector database implementations and we look forward to further partnering with the community on these endeavors.
Acknowledgements.
We thank Yury Malkov, the lead author of (Malkov and Yashunin, 2016) and maintainer of the hnswlib library, for valuable insights into the original HNSW experimental evaluation as well as thorough feedback on this manuscript.

### v2

@@ 6.5. Hub-Highway Nodes Enable Fast Traversal
where $|B_{i}|$ is the bin size (fixed to $30$ in our analysis). By averaging this value over all queries, we can plot the likelihood of visiting a hub as the search progresses.
Figure 10 shows results for the Gist, GloVe, Microsoft SpaceV and Yandex-DEEP benchmark datasets. We observe that queries tend to concentrate in the highway structures early in search, shown by the high percentage of hub nodes visited in the first 5-10% of the search steps. This result suggests that the highway allows queries to quickly navigate the similarity search graph until they find the region of the graph best suited for deep exploration. The propensity of the query to visit hubs appears to be tied to the hubness properties of the dataset. For example, the GloVe dataset uses the angular distance and has less pronounced hubs (Figure 9), and queries spend a lower percentage of their time in hub nodes in this dataset (Figure 10(b)). On the other hand, the GIST dataset has some of the highest rates of highway utilization and is also our highest-dimensional $\ell_{2}$ dataset.
@@
Figure 10. Highway nodes allow queries to traverse the graph faster. We observe that early in search,
@@ 6.6. Discussion
Small-World Graphs: The network science research community has known for decades that long-range connections and hubs induce the formation of “small-world” graphs that are easily traversed (Watts and Strogatz, 1998). This idea has been enormously influential in ANNS, providing the motivation for both NSW and HNSW (Malkov et al., 2012, 2014), but our results suggest that ANN graphs constructed over low-dimensional datasets may not in fact exhibit small-world properties. Because the $k$-occurrence distribution is near-uniform for intrinsically low-dimensional data distributions, a pure $k$NN graph (without pruning or long-range links) will not create hubs with a high in-degree. We believe that hierarchical structures are helpful in low dimensions because they help to induce hub behavior, by ensuring that search always begins from a small set of nodes. However, this is not necessary to produce hubs and induce small-world properties in high dimensions. The $k$NN graph construction process is sufficient on its own, because the hub highway emerges in high dimensions.
Scientific Implications: Our work reveals that the hub highway is an intrinsic and naturally-forming structure in high-dimensional proximity graphs. We believe this observation is both novel and has important implications for the scaling potential of many traversal heuristics. Specifically, we expect the benefits of sophisticated search initialization methods to decay under hub-style preferential attachment.
Initialization is a recurring and popular research direction for graph-based near-neighbor search, and leading algorithms vary greatly in their initialization techniques. The research question dates back to the seminal 1993 paper by Arya and Mount (Arya and Mount, 1993), which conjectured that clever search initialization – in their case, via $kd$-tree – could improve performance over random initialization. Over the following three decades, the research community has investigated diverse stratified sampling based on clusters (Sebastian and Kimia, 2002), vantage-point trees (Iwasaki, 2010), seeds formed from the graph expansion of $kd$-trees (Iwasaki and Miyazaki, 2018) and previously-visited nodes (Wang and Li, 2012), hierarchical graphs (Malkov and Yashunin, 2016), hierarchical clustering (Munoz et al., 2019), dataset medoids (Jayaram Subramanya et al., 2019), and several other methods before finally returning to cluster-stratified candidates (Oguri and Matsui, 2024; Ni et al., 2023) and random entry nodes (Jaiswal et al., 2022).
How is it that these early papers on graph search show performance gains, even as today’s best vector databases return to simple initializations without performance loss? Our hub highway hypothesis offers a clear explanation for this apparent contradiction: In the early 2000s and 2010s, datasets were low-dimensional and initialization was important to avoid local minima and long graph detours. However, modern vector databases contain data that is sufficiently high-dimensional to naturally form a fast-routing structure, explaining why initialization no longer drives performance. Based on our results, we conjecture that the largest algorithmic improvements to graph-based ANNS should come from optimizations that affect the connectivity and cost of traversal in the base graph, such as link pruning and search algorithm design.
@@ 7. Conclusion
Approximate near neighbor search has become an increasingly crucial computational workload in recent years with the seminal Hierarchical Navigable Small World (HNSW) algorithm continuing to garner significant interest and adoption from practitioners. To that end, this paper presents the first comprehensive study on the utility of the hierarchical component of HNSW over numerous large-scale datasets and performance metrics. Ultimately, we find that the hierarchy of HNSW provides no clear benefit on high-dimensional datasets and can be removed without any discernible loss in performance while providing memory savings.
@@
As part of our benchmarking study and analysis, we also release an open-source, high-performance flat navigable small world implementation, called flatnav, that we hope will serve as an additional resource to both researchers and industry practitioners. We also hope that our work can inspire future directions in algorithm design for ANN search, such as in better strategies for exploiting the hub highway structure we observed or even designing new algorithms that do manage to leverage the idea of hierarchy effectively. In the short term, we believe that our results provide immediate implications for practitioners seeking to save memory or simplify their vector database implementations and we look forward to further partnering with the community on these endeavors.